OpenStudy (anonymous):

can someone help me answer 9 questions in 30 min

OpenStudy (anonymous):

Let f(x) = 4x - 5 and g(x) = 6x - 3. Find f(x) g(x). A. 10x – 8 B. 10x – 2 C. –2x – 8 D. –2x – 2

OpenStudy (anonymous):

its not A

OpenStudy (anonymous):

What is the inverse of the given relation? y = 3x + 9

OpenStudy (sirm3d):

is the problem f(x) times g(x), or f(x) MINUS g(x), or perhaps f(x) PLUS g(x) ?

OpenStudy (anonymous):

idk what i posted is what the question says

OpenStudy (sirm3d):

looking at the choices, and if you say A is wrong, then the operation is subtraction. so my answer is (D) -2x -2 assuming the operation is subtraction

OpenStudy (sirm3d):

what are the choices for the inverse problem of \[y=3x+9\]

OpenStudy (anonymous):

y = 1/3x + 3

OpenStudy (anonymous):

A

OpenStudy (anonymous):

C. y = 3x + 3

OpenStudy (anonymous):

d. y = 1/3x - 3

OpenStudy (sirm3d):

(D) \[y=\frac{ x-9 }{ 3 }\] is the same as \[y=\frac{ 1 }{ 3 }x-3\]

OpenStudy (anonymous):

\[What is the simplest form of the radical expression?2^{4}\sqrt{2x} + 6^{4}\sqrt{2x}\]

OpenStudy (anonymous):

wth lol

OpenStudy (anonymous):

What is the simplest form of the radical expression?

OpenStudy (sirm3d):

the problem is similar to 2y + 6y

OpenStudy (anonymous):

\[b. 8^{4}\sqrt{2x}\]

OpenStudy (anonymous):

\[c. 16^{4}\sqrt{2x}\]

OpenStudy (anonymous):

D. not possible to simplify

OpenStudy (anonymous):

please show steps!!

OpenStudy (anonymous):

that are understandable

OpenStudy (anonymous):

when i first did this i added 2 and the 6 is that wrong

OpenStudy (anonymous):

i believe its b

OpenStudy (sirm3d):

\[2\sqrt[4]{2x}+6\sqrt[4]{2x}\] \[=(2+6)\sqrt[4]{2x}\] \[=8\sqrt[4]{2x}\]

OpenStudy (anonymous):

i was correct than you for checking me answer

OpenStudy (anonymous):

What is the simplest form of the quotient?

OpenStudy (anonymous):

\[^{4}\sqrt{400}\over ^{4}\sqrt{5}\]

OpenStudy (anonymous):

i got

OpenStudy (anonymous):

\[^{4}\sqrt{80}\]

OpenStudy (anonymous):

which is wrong lol

OpenStudy (sirm3d):

\[\frac{ \sqrt[4]{400} }{ \sqrt[4]{5} }\] \[=\sqrt[4]{ \frac{ 400 }{ 5 }}\] \[=\sqrt[4]{80}\]\[=\sqrt[4]{(2^4)5}\] \[2\sqrt[4]{5}\]

OpenStudy (sirm3d):

\[\sqrt[4]{80}\] can be simplified further

OpenStudy (anonymous):

haha i have no idea how i got what i got lol

OpenStudy (anonymous):

o thats right opps i forgot

OpenStudy (anonymous):

What is the solution of the equation?

OpenStudy (anonymous):

\[4(3-x)^{4/3}-5 = 59\]

OpenStudy (anonymous):

i got 5 but im thinking its -5 , 11

OpenStudy (sirm3d):

is the exponent 4/3 ?

OpenStudy (anonymous):

yes

OpenStudy (sirm3d):

\[4(3-x^{4/3})=59+5\] \[(3-x)^{4/3}=\frac{ 64 }{ 4 }\]

OpenStudy (anonymous):

16

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

i confused myself

OpenStudy (sirm3d):

\[\left( (3-x)^{4/3} \right)^{3/4}=(16)^{3/4}\]

OpenStudy (sirm3d):

oh, it's a plus or minus before (16)^(3/4)

OpenStudy (sirm3d):

\[(3-x)=\pm 8\] so x=-5 or x=11

OpenStudy (anonymous):

yay i was right :D

OpenStudy (anonymous):

Graph the equation.

OpenStudy (anonymous):

\[y = \sqrt{x} - 3\]

OpenStudy (anonymous):

my calcator doesnt give me a right answer

OpenStudy (anonymous):



OpenStudy (anonymous):

i believe this is what my answer should be

OpenStudy (sirm3d):